SPECTROSCOPY OF THE ν_{1} + ν_{2} − ν_{3}, 2ν_{1} − ν_{3}, 2ν_{1}, 3ν_{1}, 2ν_{1} + ν_{3} AND 3ν_{1} + ν_{3} OVERTONE AND COMBINATION BANDS OF BrNO

\underline{PATRICK E. GODFREY}, Air Force Research Laboratory, 3109 P Street, Wright-Patterson Air Force Base, OH 45433- 7700; GLEN P. PERRAM, Department of Engineering Physics, Air Force Institute of Technology, Wright-Patterson Air Force Base, OH 45433.

The rotation structure of the ν_{1} + ν_{2} − ν_{3}, 2ν_{1} − ν_{3}, 2ν_{1}, 3ν_{1}, 2ν_{1} + ν_{3} and 3ν_{1} + ν_{3} overtone and combination bands of BrNO were examined at high resolution (.005 - .01 cm^{−1}) using Fourier Transform Spectroscopy. Over 4900 spectral lines were recorded for the two isotopomers ^{79}BrNO and ^{81}BrNO. The rotational and quartic distortion terms, inertia defects and asymmetry parameters were determined for the (200), (300), (001), (200), (301) and (110) vibrational levels. The first-order vibration-rotation interaction terms were also determined.